OpenStudy (anonymous):

the area of the den is 3^4 square feet. the area of the kitchen is 3^3 square feet. how many times larger is the den than the kitchen?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3^4=3^3*3^1 3(kitchen area)=den area

OpenStudy (campbell_st):

really \[3^4/3^3 = 3\]